OpenStudy (anonymous):

If f(x) = x + 4, find f(3).

OpenStudy (anonymous):

7

OpenStudy (anonymous):

f(3) means x=3 y=f(3)

OpenStudy (anonymous):

one more Laura rents a movie for a flat fee of $2.00 plus an additional $0.50 for each night she keeps the movie. Choose the cost function that represents this scenario if x equals the number of nights Laura has the movie.

OpenStudy (anonymous):

those are the choices: c(x) = 2.00x + 0.50 c(x) = 2.00 + 0.50x c(x) = 2.50x c(x) = (2.00 + 0.50)x

OpenStudy (anonymous):

option two,

OpenStudy (anonymous):

\[c(x)=2+0.5x\]

OpenStudy (anonymous):

thanks